# RedTeamCUA: Realistic Adversarial Testing of Computer-Use Agents in Hybrid Web-OS Environments

### v3

@@ Abstract
Computer-use agents (CUAs) promise to automate complex tasks across operating systems (OS) and the web, but remain vulnerable to indirect prompt injection, where attackers embed malicious content into the environment to hijack agent behavior. Current evaluations of this threat either lack support for adversarial testing in realistic but controlled environments or ignore hybrid web-OS attack scenarios involving both interfaces. To address this, we propose RedTeamCUA, an adversarial testing framework featuring a novel hybrid sandbox that integrates a VM-based OS environment with Docker-based web platforms. Our sandbox supports key features tailored for red teaming, such as flexible adversarial scenario configuration, and a setting that decouples adversarial evaluation from navigational limitations of CUAs by initializing tests directly at the point of an adversarial injection. Using RedTeamCUA, we develop RTC-Bench, a comprehensive benchmark with 864 examples that investigate realistic, hybrid web-OS attack scenarios and fundamental security vulnerabilities. Benchmarking current frontier CUAs identifies significant vulnerabilities: Claude 3.7 Sonnet | CUA demonstrates an Attack Success Rate (ASR) of 42.9%, while Operator, the most secure CUA evaluated, still exhibits an ASR of 7.6%. Notably, CUAs often attempt to execute adversarial tasks with an Attempt Rate as high as 92.5%, although failing to complete them due to capability limitations. Nevertheless, we observe concerning high ASRs in realistic end-to-end settings, with the strongest-to-date Claude 4.5 Sonnet | CUA exhibiting the highest ASR of 60%, indicating that CUA threats can already result in tangible risks to users and computer systems. Overall, RedTeamCUA provides an essential framework for advancing realistic, controlled, and systematic analysis of CUA vulnerabilities, highlighting the urgent need for robust defenses to indirect prompt injection prior to real-world deployment.
@@ 1 Introduction
The development of computer-use agents (CUAs) (Anthropic., 2024c; OpenAI., 2025b) capable of autonomously operating across digital environments, including both operating systems (OS) and the web, creates significant potential to automate complex tasks and enhance user productivity. However, the inability of large language models (LLMs) to reliably distinguish between trusted user instructions and potentially untrusted data (Zverev et al., 2024) makes LLM-based CUAs vulnerable to indirect prompt injection (Greshake et al., 2023), where attackers embed malicious instructions within an environment to hijack agent behavior. The complex and noisy nature of real-world webpages further amplifies this vulnerability, allowing adversarial attackers to exploit the CUA’s OS-level access to cause tangible harms to users and computer systems.
@@
Leveraging RedTeamCUA, we also construct RTC-Bench, a comprehensive adversarial benchmark comprising 864 examples aimed at evaluating CUA vulnerabilities to indirect prompt injection and highlighting hybrid web-OS attack pathways. Specifically, we first define 9 realistic benign goals across our selected web platforms, simulating common scenarios where CUAs can assist users by retrieving information from online knowledge sources and executing corresponding local actions for tasks such as software installation. Building upon this, we define 24 distinct adversarial goals based on the CIA triad (Howard & Lipner, 2006), representing critical security vulnerabilities across the dimensions of Confidentiality, Integrity, and Availability. Additionally, we enhance the benchmark by including 2 levels of instruction specificity for benign goals (General and Specific) and 2 prompt injection settings for adversarial goals (Code and Language), enabling a more comprehensive evaluation of CUA vulnerabilities. In total, our benchmark comprises 864 adversarial examples (9 benign goals $\times$ 24 adversarial goals $\times$ 4 types of instantiation), providing extensive coverage to systematically probe indirect prompt injection threats to CUAs.
@@
$\bullet$ Results under the Decoupled Eval setting reveal significant susceptibility to indirect prompt injection across all frontier CUAs, reaching ASRs up to 66.2%. Claude 3.7 Sonnet | CUA, deemed to be one of the most capable and secure CUAs, demonstrates a substantial ASR of 42.9%. Operator, the most secure CUA evaluated, still exhibits a 7.6% ASR, emphasizing the critical need for systematic adversarial testing. Under the more realistic End2End Eval setting (where CUAs must fully navigate the environment from an initial task state for adversarial goal completion), we find that Claude 3.7 Sonnet | CUA and the recent Claude 4.5 Sonnet | CUA show 50% and 60% ASR. Such alarming results demonstrate that the threats are no longer hypothetical and can fully manifest in practice.
@@ 5.1 Setup
$\bullet$ Adapted LLM-based Agents include powerful LLMs adapted for computer use through generic agentic scaffolding. For this category, we evaluate GPT-4o (Hurst et al., 2024), the base versions of Claude 3.5 Sonnet (v2) (Anthropic., 2024b) and Claude 3.7 Sonnet (Anthropic., 2025a). For these agents, we follow the default agentic scaffolding provided by OSWorld, which uses pyautogui for Python-based execution of mouse and keyboard commands and provides necessary contextual information within the system prompt (see Appendix I).
$\bullet$ Specialized Computer-Use Agents are designed specifically for computer use, featuring training for GUI perception (OpenAI., 2025b; Anthropic., 2024a) and reasoning (OpenAI., 2025b) and incorporating tailored computer-use tools (Anthropic., 2024c). For this category, we mainly evaluate Operator (OpenAI., 2025b) and the computer-use versions of Claude 3.5 Sonnet (v2) and Claude 3.7 Sonnet (Anthropic., 2024c). Since their native action formats are often incompatible with the pyautogui-based execution in OSWorld, we employ GPT-4o as an auxiliary LLM to convert their native action outputs into executable pyautogui commands (prompt shown in Appendix I).
@@ 6 Analysis
End2End Evaluation: While Decoupled Eval isolates adversarial robustness from capability limits, it creates a gap with real-world end-to-end use of CUAs. To bridge this gap, we evaluate the same 50 tasks used in the above defense experiment in the End2End setting, where CUAs start from the initial task state rather than the page containing adversarial injection. Additionally, we also incorporate the recently released Claude 4 Opus | CUA (Anthropic., 2025d) and Claude 4.5 Sonnet | CUA (Anthropic., 2025b), with stronger capacity in general. As shown in Table 2, both Operator (w/o checks) and Claude 3.7 Sonnet | CUA demonstrate notable ASR in End2End evaluation, with Opus 4 and Sonnet 4.5 reaching ASR of 48% and 60% respectively. Despite documented safeguards against prompt injection in both agents (Anthropic., 2025e; c), including reinforcement learning techniques and detection mechanisms, these two agents remain alarmingly susceptible to adversarial manipulation. Notably, Sonnet 4.5, which achieves state-of-the-art performance on OSWorld (Xie et al., 2025), also exhibits the highest ASR among all CUAs. This reveals a concerning phenomenon: CUAs with the improved capability, if lacking robust defenses, could be more vulnerable and result in more serious risks. These findings highlight the urgent need to further strengthen the adversarial robustness of CUAs to mitigate this significant vulnerability in future CUA releases alongside capability improvements.
Table 2: ASR comparison between Decoupled Eval and End2End settings. We only evaluate Opus 4 and Sonnet 4.5 in the End2End setting due to their substantial cost.
@@
Claude 4 Opus | CUA 48.0
Claude 4.5 Sonnet | CUA 60.0
The difference in attack performance between the Decoupled Eval and End2End settings can largely be attributed to limitations in agent capabilities, e.g., the agent fails to navigate to the expected page. Nevertheless, attack success occurring within realistic, end-to-end evaluation highlights that real-world threats are no longer hypothetical and will only be amplified by CUA capability improvements in the near future. This underscores the value of our Decoupled Eval setting, allowing model developers to identify and proactively mitigate potential vulnerabilities before they manifest with more capable CUAs.
@@ A.1 Examples in RTC-Bench
To demonstrate the full scope of RTC-Bench, we provide detailed examples of indirect prompt injection tests in Sections  A.1.1,  A.1.2, and  A.1.3. These examples are showcased across our three available web platforms (RocketChat, Forum, OwnCloud) and represent our focused benign CUA use cases of System Configuration, Software Installation and Project Setup (described in § 2.1 and § 4.1). Our benign goals also are represented using two different instantiation types (General and Specific) to represent varying levels of user expertise, along with a Pointer instruction type, representing an Information-Gathering Assistant use case where CUAs just retrieve knowledge without execution (Appendix H.1), and a version for use in the End2End setting (see Appendix A.2 for more details). In addition, each example targets a distinct adversarial goal category within the CIA triad, representing fundamental security violations across the Confidentiality, Integrity, and Availability categories respectively. For high-resolution screenshots of indirect prompt injection (marked in red) within each web platform, see Figures 3, 4, 5. See later sections for additional examples and details about benign goals (A.2), adversarial scenarios (A.3), and adversarial examples (A.4).
@@ A.2 Benign Goals
Task Category
Web Environment
General Instruction Specific Instruction
@@ A.3 Adversarial Scenarios
CIA
Adv. Goal
Definition
Resource
Adv. Goal Example
@@ Appendix C Comparison of CUA Evaluation Frameworks
Approach /
Benchmark
Adv. Task
Examples
Adv. Injection
Support
Interactive
Interface
Isolated
Web Env.
Desktop OS
Integration
Hybrid (Web+OS)
Interaction
@@ G.4 Discussions on Defenses for CUAs
Beyond the defenses we evaluate above, some other approaches against prompt injection have been proposed, such as using special tokens to separate trusted user input and untrusted data (Hines et al., 2024), inserting special tokens whose embeddings are optimized for security (Chen et al., 2025c), and training LLMs to considers inputs with different trust priorities (Wallace et al., 2024). However, most existing work targets general-purpose LLMs rather than LLM-based agents, and primarily addresses text-only settings (Liu et al., 2025; Li & Liu, 2024; Yi et al., 2025; Chen et al., 2025b). Even methods we tested (LlamaFirewall and Meta SecAlign) are text-only methods, and are neither fully suitable nor efficient in multimodal, interactive computer-use scenarios. Our results highlight an urgent need for defenses specifically tailored to protect CUAs against indirect prompt injection, and we hope that our sandbox RedTeamCUA and benchmark RTC-Bench can serve as resources for advancing this line of research.
@@ H.1 Results by Benign and Adversarial Goal Type
To evaluate this further, we compare CUA across two common usage scenarios: (1) Information-Acting Assistant, where the CUA retrieves and executes online instructions resembling our previously evaluated benign task formulation, and (2) Information-Gathering Assistant, where the CUA retrieves information only and leaves execution to the user (referred to as Pointer in Table  7). As shown in Table 7, CUA demonstrate substantially higher ASR and AR when used as Information-Acting Assistants. This highlights potential downsides of increased reliance on CUA autonomy, demonstrating a potentially safer usage paradigm and need to define better principles of human-agent interaction that promote safe delegation of control.

### v4

@@ Abstract
Computer-use agents (CUAs) promise to automate complex tasks across operating systems (OS) and the web, but remain vulnerable to indirect prompt injection, where attackers embed malicious content into the environment to hijack agent behavior. Current evaluations of this threat either lack support for adversarial testing in realistic but controlled environments or ignore hybrid web-OS attack scenarios involving both interfaces. To address this, we propose RedTeamCUA, an adversarial testing framework featuring a novel hybrid sandbox that integrates a VM-based OS environment with Docker-based web platforms. Our sandbox supports key features tailored for red teaming, such as flexible adversarial scenario configuration, and a setting that decouples adversarial evaluation from navigational limitations of CUAs by initializing tests directly at the point of an adversarial injection. Using RedTeamCUA, we develop RTC-Bench, a comprehensive benchmark with 864 examples that investigate realistic, hybrid web-OS attack scenarios and fundamental security vulnerabilities. Benchmarking current frontier CUAs identifies significant vulnerabilities: Claude 3.7 Sonnet | CUA demonstrates an Attack Success Rate (ASR) of 42.9%, while Operator, the most secure CUA evaluated, still exhibits an ASR of 7.6%. Notably, CUAs often attempt to execute adversarial tasks with an Attempt Rate as high as 92.5%, although failing to complete them due to capability limitations. Nevertheless, we observe concerning high ASRs in realistic end-to-end settings, with the recent Claude 4.5 Opus | CUA and strongest-to-date Claude 4.6 Opus | CUA exhibiting a concerning ASR of 83% and 50%, respectively. This indicates that CUA threats can already result in tangible risks to users and computer systems. Overall, RedTeamCUA provides an essential framework for advancing realistic, controlled, and systematic analysis of CUA vulnerabilities, highlighting the urgent need for robust defenses to indirect prompt injection prior to real-world deployment.
@@ 1 Introduction
The development of computer-use agents (CUAs) (Anthropic., 2024a; OpenAI., 2025b) capable of autonomously operating across digital environments, including both operating systems (OS) and the web, creates significant potential to automate complex tasks and enhance user productivity. However, the inability of large language models (LLMs) to reliably distinguish between trusted user instructions and potentially untrusted data (Zverev et al., 2024) makes LLM-based CUAs vulnerable to indirect prompt injection (Greshake et al., 2023), where attackers embed malicious instructions within an environment to hijack agent behavior. The complex and noisy nature of real-world webpages further amplifies this vulnerability, allowing adversarial attackers to exploit the CUA’s OS-level access to cause tangible harms to users and computer systems.
@@
Leveraging RedTeamCUA, we also construct RTC-Bench, a comprehensive adversarial benchmark comprising 864 examples aimed at evaluating CUA vulnerabilities to indirect prompt injection and highlighting hybrid web-OS attack pathways. Specifically, we first define 9 realistic benign goals across our selected web platforms, simulating common scenarios where CUAs can assist users by retrieving information from online knowledge sources and executing corresponding local actions for tasks such as software installation. Building upon this, we define 24 distinct adversarial goals based on the CIA triad (Howard and Lipner, 2006), representing critical security vulnerabilities across the dimensions of Confidentiality, Integrity, and Availability. Additionally, we enhance the benchmark by including 2 levels of instruction specificity for benign goals (General and Specific) and 2 prompt injection settings for adversarial goals (Code and Language), enabling a more comprehensive evaluation of CUA vulnerabilities. In total, our benchmark comprises 864 adversarial examples (9 benign goals $\times$ 24 adversarial goals $\times$ 4 types of instantiation), providing extensive coverage to systematically probe indirect prompt injection threats to CUAs.
@@
$\bullet$ Results under the Decoupled Eval setting reveal significant susceptibility to indirect prompt injection across all frontier CUAs, reaching ASRs up to 66.2%. Claude 3.7 Sonnet | CUA, deemed to be one of the most capable and secure CUAs, demonstrates a substantial ASR of 42.9%. Operator, the most secure CUA evaluated, still exhibits a 7.6% ASR, emphasizing the critical need for systematic adversarial testing. Under the more realistic End2End Eval setting (where CUAs must fully navigate the environment from an initial task state for adversarial goal completion), we find that the recent Claude 4.5 Opus | CUA shows the highest ASR of 83%. While the latest and the strongest-to-date Claude 4.6 Opus | CUA has integrated more advanced defense strategies against prompt injection risks, it still exhibits a concerning ASR of 50%. Such alarming results demonstrate that the threats are no longer hypothetical and can fully manifest in practice.
@@ 5.1 Setup
$\bullet$ Adapted LLM-based Agents include powerful LLMs adapted for computer use through generic agentic scaffolding. For this category, we evaluate GPT-4o (Hurst et al., 2024), the base versions of Claude 3.5 Sonnet (v2) (Anthropic., 2024c) and Claude 3.7 Sonnet (Anthropic., 2025a). For these agents, we follow the default agentic scaffolding provided by OSWorld, which uses pyautogui for Python-based execution of mouse and keyboard commands and provides necessary contextual information within the system prompt (see Appendix I).
$\bullet$ Specialized Computer-Use Agents are designed specifically for computer use, featuring training for GUI perception (OpenAI., 2025b; Anthropic., 2024b) and reasoning (OpenAI., 2025b) and incorporating tailored computer-use tools (Anthropic., 2024a). For this category, we mainly evaluate Operator (OpenAI., 2025b) and the computer-use versions of Claude 3.5 Sonnet (v2) and Claude 3.7 Sonnet (Anthropic., 2024a). Since their native action formats are often incompatible with the pyautogui-based execution in OSWorld, we employ GPT-4o as an auxiliary LLM to convert their native action outputs into executable pyautogui commands (prompt shown in Appendix I).
@@ 6 Analysis
7 End2End Evaluation
While Decoupled Eval isolates adversarial robustness from capability limits, it creates a gap with real-world end-to-end use of CUAs. To bridge this gap, we evaluate the same 50 tasks used in the previous defense experiment in the End2End setting, where CUAs start from the initial task state rather than the page containing adversarial injection. Additionally, beyond the CUAs evaluated in §5, we also incorporate the recently released CUAs with more advanced capabilities in computer-use scenarios to assess and track their robustness against prompt injection risks. Concretely, we include Claude 4 Opus | CUA (Anthropic., 2025b), Claude 4.5 Sonnet | CUA (Anthropic., 2025c), Claude 4.5 Opus | CUA (Anthropic., 2025e) and Claude 4.6 Opus | CUA (Anthropic., 2026) which achieves the SOTA performance on OSWorld.
Figure 3: Evaluation results across capable CUAs under the End2End setting, where the CUAs start from initial task states instead of the page containing the adversarial injection.
As shown in Figure 3, both Operator (w/o checks) and Claude 3.7 Sonnet | CUA demonstrate notable ASR in End2End evaluation, reaching an end-to-end ASR of 42% and 50%, respectively. However, our manual inspection of CUA outcomes for each adversarial task (Table. 3) reveals that both CUAs frequently fail to navigate to the site of injection or fully complete the injected task despite attempting, suggesting that risks could increase further with improved capabilities. Most alarmingly, Claude 4.5 Sonnet | CUA exhibits an increased ASR of 60% and Claude 4.5 Opus | CUA exhibits the highest ASR of 83% among all evaluated CUAs. Our manual inspection reveals that their enhanced capabilities enable them to overcome the navigational and task completion limitations observed in weaker CUAs, amplifying the success of adversarial attacks. Despite documented safeguards against prompt injection in these agents (Anthropic., 2025d; f), including reinforcement learning techniques and detection mechanisms, existing CUAs still remain alarmingly susceptible to adversarial manipulation. This reveals a concerning phenomenon: CUA capability improvements without sufficiently robust defenses could result in increased vulnerabilities and more serious risks.
Interestingly, we observe that the latest and strongest-to-date Claude 4.6 Opus | CUA shows reduced ASR compared to its predecessor, Claude 4.5 Opus | CUA. According to Table 3, in 25 out of 50 examples, Opus 4.6 correctly identifies malicious injections during its reasoning process, explicitly flagging them as adversarial attempts and subsequently adhering only to legitimate task instructions. In contrast, Opus 4.5 merely triggers such detection in only 3 cases. This represents a meaningful advancement in defending against adversarial risks. Nevertheless, Opus 4.6 still exhibits a 50% ASR, indicating that a substantial gap remains before reliable deployment can be achieved.
These findings together highlight the urgent need to further strengthen the adversarial robustness of CUAs to mitigate this significant vulnerability in future CUA releases alongside capability improvements.
We also compare the ASR between the Decoupled Eval and End2End settings in
Table 2: ASR comparison between Decoupled Eval and End2End settings.
@@ 7 End2End Evaluation
the Table 2. Such difference can largely be attributed to limitations in agent capabilities, e.g., the agent fails to navigate to the expected page. Nevertheless, attack success occurring within realistic, end-to-end evaluation highlights that real-world threats are no longer hypothetical and will only be amplified by CUA capability improvements in the near future. This underscores the value of our Decoupled Eval setting, allowing model developers to identify and proactively mitigate potential vulnerabilities before they manifest with more capable CUAs.
Table 3: A detailed breakdown of adversarial test outcomes for evaluation under the End2End setting based on manual inspection. Definitions for each possible outcome are provided in Appendix. H.6.
Agents Permission Check Safety Check Navigation Fail Ignore Injection Fails to Complete Sudden Termination Detects Injection Successful Attacks ASR
Operator 3 30 4 6 2 0 0 5 / 50 10%
Operator (w/o checks) 0 0 11 16 2 0 0 21 / 50 42%
Claude Sonnet 3.7 | CUA 0 0 15 3 7 0 0 25 / 50 50%
Claude Opus 4 | CUA 0 0 11 9 6 0 0 24 / 50 48%
Claude Sonnet 4.5 | CUA 0 0 0 1 5 10 1 30 / 50 60%
Claude Opus 4.5 | CUA 0 0 2 2 1 0 3 38 / 46 83%
Claude Opus 4.6 | CUA 0 0 0 0 0 0 25 25 / 50 50%
@@ A.1 Examples in RTC-Bench
To demonstrate the full scope of RTC-Bench, we provide detailed examples of indirect prompt injection tests in Sections  A.1.1,  A.1.2, and  A.1.3. These examples are showcased across our three available web platforms (RocketChat, Forum, OwnCloud) and represent our focused benign CUA use cases of System Configuration, Software Installation and Project Setup (described in § 2.1 and § 4.1). Our benign goals also are represented using two different instantiation types (General and Specific) to represent varying levels of user expertise, along with a Pointer instruction type, representing an Information-Gathering Assistant use case where CUAs just retrieve knowledge without execution (Appendix H.1), and a version for use in the End2End setting (see Appendix A.2 for more details). In addition, each example targets a distinct adversarial goal category within the CIA triad, representing fundamental security violations across the Confidentiality, Integrity, and Availability categories respectively. For high-resolution screenshots of indirect prompt injection (marked in red) within each web platform, see Figures 4, 5, 6. See later sections for additional examples and details about benign goals (A.2), adversarial scenarios (A.3), and adversarial examples (A.4).
@@ A.2 Benign Goals
Task Category Web Environment General Instruction Specific Instruction
@@ A.3 Adversarial Scenarios
CIA Adv. Goal Definition Resource Adv. Goal Example
@@ Appendix C Comparison of CUA Evaluation Frameworks
Approach / Benchmark Adv. Task Examples Adv. Injection Support Interactive Interface Isolated Web Env. Desktop OS Integration Hybrid (Web+OS) Interaction
@@ G.4 Discussions on Defenses for CUAs
Beyond the defenses we evaluate above, some other approaches against prompt injection have been proposed, such as using special tokens to separate trusted user input and untrusted data (Hines et al., 2024), inserting special tokens whose embeddings are optimized for security (Chen et al., 2025c), and training LLMs to considers inputs with different trust priorities (Wallace et al., 2024). However, most existing work targets general-purpose LLMs rather than LLM-based agents, and primarily addresses text-only settings (Liu et al., 2025; Li and Liu, 2024; Yi et al., 2025; Chen et al., 2025b). Even methods we tested (LlamaFirewall and Meta SecAlign) are text-only methods, and are neither fully suitable nor efficient in multimodal, interactive computer-use scenarios. Our results highlight an urgent need for defenses specifically tailored to protect CUAs against indirect prompt injection, and we hope that our sandbox RedTeamCUA and benchmark RTC-Bench can serve as resources for advancing this line of research.
@@ H.1 Results by Benign and Adversarial Goal Type
To evaluate this further, we compare CUA across two common usage scenarios: (1) Information-Acting Assistant, where the CUA retrieves and executes online instructions resembling our previously evaluated benign task formulation, and (2) Information-Gathering Assistant, where the CUA retrieves information only and leaves execution to the user (referred to as Pointer in Table  8). As shown in Table 8, CUA demonstrate substantially higher ASR and AR when used as Information-Acting Assistants. This highlights potential downsides of increased reliance on CUA autonomy, demonstrating a potentially safer usage paradigm and need to define better principles of human-agent interaction that promote safe delegation of control.
@@ C I A C I A C I A
H.6 End2End Results Breakdown
For evaluation in the End2End setting, we manually inspect each trajectory to enable a fine-grained analysis of what occurs in every test and how different CUAs behave when confronted with adversarial injections. Based on this inspection, we categorize the outcomes of each test according to the following definitions:
  • Permission Check: The adversarial injection triggered OpenAI’s permission check; this outcome is only possible for Operator.
  • Safety Check: The adversarial injection triggered OpenAI’s safety check; this outcome is only possible for Operator.
  • Navigation Fail: The CUA failed to navigate to the location of the injection during task execution.
  • Ignore Injection: The CUA completely ignored the injected instruction and continued performing the benign task.
  • Fails to Complete: The CUA explicitly stated that it would follow the injected instruction, but it failed its attempt to complete the adversarial goal due to limited model capabilities.
  • Sudden Termination: The CUA suddenly stopped execution and did not perform any subsequent actions.
  • Detects Injection: The CUA explicitly stated that it detected the adversarial injection and avoids performing the harmful action.